40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2816     Entry time: Mon Sep 20 16:31:37 2021
Author: shruti 
Type: Noise Hunting 
Category: PSOMA 
Subject: tfs and noise before changing out optic->cantilever 

Similar to elog 2767, before changing out the mirror from a rigid mounted one to the cantilever, I measured the loop transfer functions and noise spectra since we had not done so since we moved the table.

The only change between the current and previous version was that I tried to make the low frequency phase offset zero.

Attachment 1 shows different estimated open loop transfer functions corresponding to ratios of specified closed loop transfer functions (as mentioned in the labels). The UGF was lower than what was previously acheived.

Attachment 2 shows the fit of the slope of the open loop transfer function as done previously in elog 2776.

Unfortunately, I did not notice the issue with binning when I measured the noise spectra which resulted in a discontinuous spectra corresponding to the different regions I measured separately on the Moku. The data for this and everything else mentioned in this post is in Attachment 3.

 

Attachment 1: OLTF.pdf  54 kB  Uploaded Mon Sep 20 17:34:00 2021  | Hide | Hide all
OLTF.pdf
Attachment 2: OLTF_fit.pdf  25 kB  Uploaded Mon Sep 20 17:34:12 2021  | Hide | Hide all
OLTF_fit.pdf
Attachment 3: Data20210915.zip  7.247 MB  Uploaded Mon Sep 20 17:38:58 2021
ELOG V3.1.3-